# CoMoE: Contrastive Representation for Mixture-of-Experts in Parameter-Efficient Fine-tuning

### v1

@@ 1 Introduction
Parameter-Efficient Fine-Tuning (PEFT) has emerged to efficiently adapt Large Language Models (LLMs) to downstream tasks by updating only a subset of parameters, significantly reducing computational and memory overhead Hu et al. (2021); Liu et al. (2022); He et al. (2021). However, it struggles with substantially increased dataset sizes, especially heterogeneous training datasets, which poses a significant practical challenge Huang et al. (2024); Wang et al. (2024). Mixture-of-Experts (MoE) offers a versatile solution to the challenge for its modular design Zhang et al. (2024).
@@
Thus, Low-rank Adaptation (LoRA), as a popular and effective PEFT method, has been widely integrated with MoE Dou et al. (2023); Li et al. (2024), leveraging MoE’s modularity to enhance the model’s capacity and performance. By sparsely activating a subset of experts, LoRA’s MoE variants achieve efficient training on heterogeneous datasets and allocate the experts adaptively Tian et al. (2024). Specifically, the sparse activation is controlled through a router mechanism (e.g., top-k routing) that dispatches inputs to the activated experts. Basically, given an input token, only a subset of specialized experts contribute to the output, while other irrelevant experts remain inactive.
@@
Ideally, each expert should specialize in distinct representation subspaces and semantic skills, thereby collaboratively enhancing the model’s representational capacity and enabling a broader spectrum of knowledge Liu et al. (2023). However, despite the explicit division into multiple experts in MoE architecture, its modularization degree remains questionable. Two issues persist: (1) expert knowledge redundancy, where insufficient specialization constraints lead to overlapping functionalities among experts, limiting the model’s capacity Feng et al. (2025); (2) expert load imbalance, where inadequate modularity and specialization during training result in frequent activation of only a subset of experts, which underutilizes other experts and contradicts its original design intent. Consequently, as some studies have indicated Qian et al. (2024), simply stacking more experts does not linearly improve performance; instead, it leads to a performance bottleneck. Existing studies propose load balance loss Li et al. (2024) and localized balancing constraint Dou et al. (2023) to alleviate the mentioned issues, but that is still far from enough.
@@
Figure 1: Given an input token $x$, (a) illustrates the workflows of top-2 routing, which serves as a fundamental mechanism of CoMoE; (b) illustrates the motivation of CoMoE: maximizing MI between input $x$ and activated experts while minimizing MI between input $x$ and inactive experts.
@@
In this paper, we propose a novel perspective to promote the specialization of experts. As illustrated in Fig. 1, building upon top-k routing, we categorize the experts into activated experts and inactivated experts. Then, we quantify the specialization of experts by mutual information (MI) between the input token and the two types of experts. To promote expert specialization, we define an MI gap, which is derived from the aforementioned MI, and aim to maximize it. In practice, based on the InfoNCE theory Oord et al. (2018), such an MI gap can be approximated via a contrastive objective by using positive samples from the activated experts and negative samples from the inactive experts Lan et al. (2024); Wen et al. (2024). The contrastive objective is incorporated as an auxiliary objective during training, encouraging specialization and modularization among experts. We name the proposed method Contrastive Representation for MoE (CoMoE), a novel MoE variant. Empirically, we evaluate CoMoE on diverse benchmarks, showcasing its remarkable performance on heterogeneous tasks. Summary of our contributions:
@@ 2 Preliminaries
LoRA Basics LoRA Hu et al. (2021) introduces a pair of low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ to reparameterize the pretrained weights $\mathbf{W_{0}}$ in a residual manner. Specifically, input $x$ is processed through both the frozen weights and the low-rank matrices:
@@ 3.1 Parameter-Efficient Fine-Tuning
PEFT methods can be categorized into Adapter Tuning, Prompt Tuning, and Prefix Tuning: Adapter Tuning Hu et al. (2023, 2021); Zhang et al. (2023b) includes algorithms like BitFit Zaken et al. (2021), which links learnable vectors to hidden states using multiplication or addition. (IA)3 Liu et al. (2022) introduces interpretable adapters to enhance model transparency and task-specific adaptability. Prompt-Tuning Li and Liang (2021); Liu et al. (2021); Zhu et al. (2024) introduces learnable prefixes to the input sequence. Prefix-Tuning Li and Liang (2021) modifies hidden states by adding embeddings, making it efficient for few-shot tasks.
LoRA Hu et al. (2021) reduces the number of tunable parameters by introducing low-rank matrices into pre-trained weights, enhancing performance and efficiency. Currently, LoRA becomes the most popular and commonly used PEFT method Cui et al. (2023); Zhang et al. (2022). Numerous studies have focused on improving LoRA: Tied-LoRA Renduchintala et al. (2023) further reduces trainable parameters by applying weight binding. AdaLoRA Zhang et al. (2023a) utilizes singular value decomposition to decompose the weight matrices and eliminates insignificant singular values, thereby simplifying the update process. DoRA Liu et al. (2024) decomposes weights into magnitude and direction components for efficiency.
@@ 3.2 Mixture of Experts
Recent research has increased its focus on combining LoRA with MoE Shazeer et al. (2017); Jacobs et al. (1991) for adaptable and scalable LLMs, emerging as a balance between model capacity and computation overhead Li et al. (2024); Zhang et al. (2024). Generally, MoE is employed for two primary purposes: (1)Avoiding catastrophic forgetting: LoRAMoE Dou et al. (2023) leverages token-based routing to mitigate knowledge loss. MOELoRA Liu et al. (2023) enhances expert selection in multi-tasking. MoRAL Yang et al. (2024) adapts to new tasks in lifelong learning while preserving old knowledge. (2) Model efficiency: HydraLoRA Tian et al. (2024) adopts an asymmetric LoRA structure without domain expertise, and MiLoRA Zhang et al. (2024) employs a prompt-aware routing to reduce latency. LoRAMoE Dou et al. (2023) incorporates a localized balancing constraint to achieve balanced workloads. SCMoE Shi et al. (2024) leverages unchosen experts to enhance parameter efficiency. Current methods focus on architectural expert partitioning but neglect capacity underutilization caused by expert redundancy.
@@ 4 Methods
[…]
@@ 4.1 Motivation
As demonstrated in OMoE Feng et al. (2025), the vanilla MoE variant lacks specialization and modularity, causing LoRA experts to collapse into similar distributions. The lack of specialization and redundant knowledge minimizes the utilization of capacity, which exacerbates performance degradation in heterogeneous tasks. Existing MoE variants Li et al. (2024); Liu et al. (2023); Luo et al. (2024) leverage balance loss to promote specialization among the experts, but fall far short.
Ideally, experts should exhibit modularity and high specialization with minimal redundancy. To quantify these properties, we leverage MI, a basic concept in information theory, to evaluate the dependence between inputs and experts in MoE. Thus, we formalize the aforementioned idea using MI to quantify the specialization and redundancy between the experts in top-k routing:
  • Maximizing the MI between inputs and activated experts: Promote activated experts to respond to inputs that highly match, thereby encouraging specialization. In addition, MI serves as an information bottleneck that filters irrelevant noise.
@@ 4.2 MI Gap for Top-k Routing
MI Gap for activated experts and inactivated tas experts in top-k routing is defined by:
@@ 4.3 Contrastive Representation for MI Gap
^$\hat{\mathcal{L}}_{\mathrm{NCE}}$ $=-\mathbb{E}_{p\left(x,e^{+}\right)}\mathbb{E}_{\mathcal{D}_{\neg\text{top-k}}}$ (10)
@@ 5.1 Experimental Settings
Datasets and Benchmarks. We conduct experiments on a collection of tasks: (a) Diverse commonsense reasoning datasets: ARC-e and ARC-c Clark et al. (2018), OpenBookQA (OBQA) Mihaylov et al. (2018), PIQA Bisk et al. (2020), SocialIQA (SIQA) Sap et al. (2019), and BoolQ Clark et al. (2019). (b) A science completion task: Hellaswag Zellers et al. (2019). (c) A fill-in-the-blank task: Winogrande Sakaguchi et al. (2021).
We utilize the PEFT framework provided by Hu et al. (2023); Li et al. (2024) for training on these datasets. We choose LLaMA-2 7B and Gemma 2B as our backbone models. The detailed statistics and evaluation metrics can be found in Appendix C.
Baselines. In this study, we compare our method with several popular and well-established baselines to assess its performance. For multi-task settings, we evaluate CoMoE against LoRA, its variants, and MoE-based methods, including: (1) LoRA Hu et al. (2021); (2) DoRA Liu et al. (2024), which decomposes LoRA weights into magnitude and direction; (3) MoELoRA Liu et al. (2023), which decomposes a LoRA module into a mixture of experts; (4) MiLoRA Zhang et al. (2024), which treats each LoRA module as an expert and employs a routing mechanism; (5) MixLoRA Li et al. (2024), a resource-efficient sparse MoE model based on LoRA; (6) OMoE Feng et al. (2025), which applies hard constraints to promote diversity. For single-task settings, we compare our approach with other PEFT baselines, including: AdaLoRA Zhang et al. (2023a); Parallel-Adapter He et al. (2021); Learned-Adapter Zhang et al. (2023b); P-tuning v2 Liu et al. (2021); IAPT Zhu et al. (2024); BitFit Zaken et al. (2021); (IA)3 Liu et al. (2022); SSP Hu et al. (2022). Most of the results are directly extracted from  Zhang et al. (2024), with a few baselines reproduced by running the provided source code. In single-task settings, we compare CoMoE with LoRA, DoRA, and MixLoRA as representative baselines, while comparisons with other methods are presented in Appendix E.
@@ 5.2 Main Results
Multi-task Setup. Table 1 summarizes the multi-task performance of CoMoE and baselines on LLaMA-2 7B. The results verify that both LoRA and DoRA, due to their lack of MoE structures, struggle to handle heterogeneous datasets, observing a significant drop in multi-task settings (7% $\sim$ 12% degradation). MoE-based PEFT methods (MoELoRA, MiLoRA, MixLoRA, and OMoE) mitigate the performance degradation in multi-task settings but fail to leverage the modularity and specialization in MoE, leaving potential performance gains unexploited. In contrast, CoMoE introduces contrastive learning to promote the modularization and specialization of experts, enabling efficient utilization of the MoE’s capacity. Thus, CoMoE not only improves parameter efficiency (reduces the number of experts) but also achieves an average accuracy gain of +1.3.
Single-task Setup. In this setup, we compare the performance of CoMoE and baselines in single-task settings. The experimental results are shown in Table 2. Comparisons with other baselines are provided in Appendix E. CoMoE demonstrates superior parameter efficiency while maintaining comparable performance, with a reduction of approximately $50\%$ in tunable parameters. Remarkably, the diversity among experts even yields accuracy improvements on a subset of datasets (e.g., ARC-e, BoolQ, and SIQA).
@@ 6 Conclusion
We show that our proposed method can significantly improve the specialization of experts and the performance of MoE in multi-task settings. However, there are several limitations to acknowledge: (a) The computational cost of CoMoE. CoMoE leverages inactivated experts as negative samples, and its computational cost is proportional to $O(N)$. However, it substantially improves expert specialization and modularization, achieving remarkable performance with a limited number of experts; (b) Other evaluation benchmarks were not considered. Nevertheless, CoMoE is simple and can be easily integrated with different backbone models and various downstream tasks.
@@ Appendix A Theoretical Proof
Proof. As illustrated in Eq. 8, we utilize $h_{1}(x,e^{+})\propto p(e^{+}|x)/p(e^{+})$ and $h_{2}(x,e^{-})\propto p(e^{-}|x)/p(e^{-})$ to approximate the information density ratio, which preserves the MI between $x$ and $M$. Based on the standard derivations presented in InfoNCE Oord et al. (2018) and Wen et al. (2024), the contrastive objective can be rewritten as:

### v2

@@ 1 Introduction
Parameter-Efficient Fine-Tuning (PEFT) has emerged to efficiently adapt Large Language Models (LLMs) to downstream tasks by updating only a subset of parameters, significantly reducing computational and memory overhead hu2022lora; liu2022few; he2021towards. However, it struggles with substantially increased dataset sizes, especially heterogeneous training datasets, which poses a significant practical challenge huang2024c; wang2024ts. Mixture-of-Experts (MoE) offers a versatile solution to the challenge for its modular design zhang2024milora.
@@
Thus, Low-rank Adaptation (LoRA), as a popular and effective PEFT method, has been widely integrated with MoE dou2024loramoe; li2024mixlora, leveraging MoE’s modularity to enhance the model’s capacity and performance. By sparsely activating a subset of experts, LoRA’s MoE variants achieve efficient training on heterogeneous datasets and allocate the experts adaptively tian2024hydralora. Specifically, the sparse activation is controlled through a router mechanism (e.g., top-k routing) that dispatches inputs to the activated experts. Basically, given an input token, only a subset of specialized experts contribute to the output, while other irrelevant experts remain inactive.
@@
Ideally, each expert should specialize in distinct representation subspaces and semantic skills, thereby collaboratively enhancing the model’s representational capacity and enabling a broader spectrum of knowledge liu2023moelora. However, despite the explicit division into multiple experts in MoE architecture, its modularization degree remains questionable. Two issues persist: (1) expert knowledge redundancy, where insufficient specialization constraints lead to overlapping functionalities among experts, limiting the model’s capacity feng2025omoe; (2) expert load imbalance, where inadequate modularity and specialization during training result in frequent activation of only a subset of experts, which underutilizes other experts and contradicts its original design intent. Consequently, as some studies have indicated qian2024scaling, simply stacking more experts does not linearly improve performance; instead, it leads to a performance bottleneck. Existing studies propose load balance loss li2024mixlora and localized balancing constraint dou2024loramoe to alleviate the mentioned issues, but that is still far from enough.
@@
Figure 1: Given an input token $x$, (a) illustrates a workflow of top-2 routing, which serves as a fundamental mechanism of CoMoE; (b) illustrates the motivation of CoMoE: maximizing MI between input $x$ and activated experts while minimizing MI between input $x$ and inactive experts.
@@
In this paper, we propose a novel perspective to promote the specialization of experts. As illustrated in Fig. 1, building upon top-k routing, we categorize the experts into activated experts and inactivated experts. Then, we quantify the specialization of experts by mutual information (MI) between the input token and the two types of experts. To promote expert specialization, we define an MI gap, which is derived from the aforementioned MI, and aim to maximize it. In practice, based on the InfoNCE theory oord2018representation, such an MI gap can be approximated via a contrastive objective by using positive samples from the activated experts and negative samples from the inactive experts lan2024contrastive; wen2024contrastive. The contrastive objective is incorporated as an auxiliary objective during training, encouraging specialization and modularization among experts. We name the proposed method Contrastive Representation for MoE (CoMoE), a novel MoE variant. Empirically, we evaluate CoMoE on diverse benchmarks, showcasing its remarkable performance on heterogeneous tasks. Summary of our contributions:
@@ 2 Preliminaries
LoRA Basics LoRA hu2022lora introduces a pair of low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ to reparameterize the pretrained weights $\mathbf{W_{0}}$ in a residual manner. Specifically, input $x$ is processed through both the frozen weights and the low-rank matrices:
@@ 3.1 Parameter-Efficient Fine-Tuning
PEFT methods can be categorized into Adapter Tuning, Prompt Tuning, and Prefix Tuning: Adapter Tuning hu2023llm; hu2022lora; zhang2023learned includes algorithms like BitFit ben2022bitfit, which links learnable vectors to hidden states using multiplication or addition. (IA)3 liu2022few introduces interpretable adapters to enhance model transparency and task-specific adaptability. Prompt-Tuning li2021prefix; liu2021p; zhu2024iapt introduces learnable prefixes to the input sequence. Prefix-Tuning li2021prefix modifies hidden states by adding embeddings, making it efficient for few-shot tasks.
LoRA hu2022lora reduces the number of tunable parameters by introducing low-rank matrices into pre-trained weights, enhancing performance and efficiency. Currently, LoRA becomes the most popular and commonly used PEFT method cui2023ultrafeedback; zhang2022pcee. Numerous studies have focused on improving LoRA: Tied-LoRA renduchintala2024tied further reduces trainable parameters by applying weight binding. AdaLoRA zhang2023adaptive utilizes singular value decomposition to decompose the weight matrices and eliminates insignificant singular values, thereby simplifying the update process. LoraHub huang2023lorahub employs dynamic LoRA composition with gradient-free weight optimization. MoDULA ma2024modula uses a mixture of domain-specific and universal LoRA with dynamic weight balancing. DoRA liu2024dora decomposes weights into magnitude and direction components for efficiency.
@@ 3.2 Mixture of Experts
Recent research has increased its focus on combining LoRA with MoE shazeer2017outrageously; jacobs1991adaptive for adaptable and scalable LLMs, emerging as a balance between model capacity and computation overhead li2024mixlora; zhang2024milora. Generally, MoE is employed for two primary purposes: (1)Avoiding catastrophic forgetting: LoRAMoE dou2024loramoe leverages token-based routing to mitigate knowledge loss. MOELoRA liu2023moelora enhances expert selection in multi-tasking. MoRAL yang2024moral adapts to new tasks in lifelong learning while preserving old knowledge. (2) Model efficiency: HydraLoRA tian2024hydralora adopts an asymmetric LoRA structure without domain expertise, and MiLoRA zhang2024milora employs a prompt-aware routing to reduce latency. LoRAMoE dou2024loramoe incorporates a localized balancing constraint to achieve balanced workloads. SCMoE shi2024unchosen leverages unchosen experts to enhance parameter efficiency. MoLoRA zadouri2023pushing achieves highly parameter-efficient MoE through sparse expert activation, low-rank decomposition, and parameter sharing. C-Poly wang2024customizable enables customizable integration of heterogeneous parameter-efficient modules through meta-learned strategies. Current methods focus on architectural expert partitioning but neglect capacity underutilization caused by expert redundancy.
@@ 4 Methods
Figure 2: Architecture of CoMoE. Activated experts $e_{i}^{t},\,i\in\text{top2}$ in top-2 routing are selected as the query $q$ and positive keys $k^{+}$, while inactivated experts $e_{i}^{t},\,i\notin\text{top2}$ serve as negative keys $k^{-}$.
@@ 4.1 Motivation
As demonstrated in OMoE feng2025omoe, the vanilla MoE variant lacks specialization and modularity, causing LoRA experts to collapse into similar distributions. The lack of specialization and redundant knowledge minimizes the utilization of capacity, which exacerbates performance degradation in heterogeneous tasks. Existing MoE variants li2024mixlora; liu2023moelora; luo2024moelora leverage balance loss to promote specialization among the experts, but fall far short.
Ideally, experts should exhibit modularity and high specialization with minimal redundancy. To quantify these properties, we leverage MI, a basic concept in information theory, to evaluate the dependence between inputs and experts in MoE. An intuitive illustration of expert specialization is provided in Appendix K. Thus, we formalize the aforementioned idea using MI to quantify the specialization and redundancy between the experts in top-k routing:
  • Maximizing the MI between inputs and activated experts: Promote activated experts to respond to inputs that highly match, thereby encouraging specialization. In addition, MI can serve as an information bottleneck that filters irrelevant noise.
@@ 4.2 MI Gap for Top-k Routing
MI Gap for activated experts and inactivated experts in top-k routing is defined by:
@@ 4.3 Contrastive Representation for MI Gap
$\hat{\mathcal{L}}_{\mathrm{NCE}}$ $=-\mathbb{E}_{p\left(x,e^{+}\right)}\mathbb{E}_{\mathcal{D}_{\neg\text{top-k}}}$ (10)
@@ 5 Experiments
Method GSM8K AddSub AQuA MultiArith SingleEQ SVAMP Arithmetic Avg.
LoRA 31.1 84.8 17.6 88.3 90.2 65.0 2.69 54.2
CoMoE-LoRA 32.2 89.9 31.4 91.7 90.2 66.0 80.9 68.90
Table 3: Performance comparison between LoRA and CoMoE-LoRA on math reasoning tasks. CoMoE-LoRA achieves substantial improvements under multi-task settings.
@@ 5.1 Experimental Settings
Datasets and Benchmarks. We conduct experiments on a collection of tasks: (a) Diverse commonsense reasoning datasets: ARC-e and ARC-c clark2018think, OpenBookQA (OBQA) mihaylov2018can, PIQA bisk2020piqa, SocialIQA (SIQA) sap2019social, and BoolQ clark2019boolq. (b) A science completion task: Hellaswag zellers2019hellaswag. (c) A fill-in-the-blank task: Winogrande sakaguchi2021winogrande. (d) Math reasoning tasks: GSM8K cobbe2021training (grade-school word problems), AddSub hosseini2014learning (single-step arithmetic), AQuA ling2017program (algebra multiple-choice), MultiArith roy2015solving (multi-step arithmetic), SingleEQ koncel2015parsing (single-equation problems), and SVAMP patel2021nlp (robustness testing). (e) Symbolic math expression evaluation: Arithmetic brown2020language.
We utilize the PEFT framework provided by hu2023llm; li2024mixlora for training on these datasets. We choose LLaMA-2 7B and Gemma 2B as our backbone models. The detailed statistics and evaluation metrics can be found in Appendix C.
Baselines. In this study, we compare our method with several popular and well-established baselines to assess its performance. For multi-task settings, we evaluate CoMoE against LoRA, its variants, and MoE-based methods, including: (1) LoRA hu2022lora; (2) DoRA liu2024dora, which decomposes LoRA weights into magnitude and direction; (3) MoELoRA liu2023moelora, which decomposes a LoRA module into a mixture of experts; (4) MiLoRA zhang2024milora, which treats each LoRA module as an expert and employs a routing mechanism; (5) MixLoRA li2024mixlora, a resource-efficient sparse MoE model based on LoRA; (6) OMoE feng2025omoe, which applies hard constraints to promote diversity. For single-task settings, we compare our approach with other PEFT baselines, including: AdaLoRA zhang2023adaptive; Parallel-Adapter he2021towards; Learned-Adapter zhang2023learned; P-tuning v2 liu2021p; IAPT zhu2024iapt; BitFit ben2022bitfit; (IA)3 liu2022few; SSP hu2022sparse.
Most of the results are directly extracted from  zhang2024milora, with a few baselines reproduced by running the provided source code. In single-task settings, we compare CoMoE with LoRA, DoRA, and MixLoRA as representative baselines, while comparisons with other methods are presented in Appendix E.
@@ 5.2 Main Results
Multi-task Setup. Table 1 summarizes the multi-task performance of CoMoE and baselines on LLaMA-2 7B. The results verify that both LoRA and DoRA, due to their lack of MoE structures, struggle to handle heterogeneous datasets, observing a significant drop in multi-task settings (7% $\sim$ 12% degradation). MoE-based PEFT methods (MoELoRA, MiLoRA, MixLoRA, and OMoE) mitigate the performance degradation in multi-task settings but fail to leverage the modularity and specialization in MoE, leaving potential performance gains unexploited. In contrast, CoMoE introduces contrastive learning to promote the modularization and specialization of experts, enabling efficient utilization of the MoE’s capacity. Thus, CoMoE not only improves parameter efficiency (reduces the number of experts) but also achieves an average accuracy gain of +1.3. In addition, we evaluate CoMoE-LoRA on several math reasoning tasks under more complex multi-task settings, as shown in Table 3. The results related to LoRA are reproduced from luo2024moelora. Compared with LoRA, CoMoE achieves a notable improvement in performance, further demonstrating its generalization in math reasoning tasks.
Single-task Setup. In this setup, we compare the performance of CoMoE and baselines in single-task settings. The experimental results are shown in Table 2. Comparisons with other baselines are provided in Appendix E. CoMoE demonstrates superior parameter efficiency while maintaining comparable performance, with a reduction of approximately $50\%$ in tunable parameters. Remarkably, the diversity among experts even yields accuracy improvements on a subset of datasets (e.g., ARC-e, BoolQ, and SIQA). The performance differences between multi-task and single-task settings are further analyzed in Appendix J.
@@ 6 Conclusion
We show that our proposed method can significantly improve the specialization of experts and the performance of MoE in multi-task settings. However, there are several limitations to acknowledge: (a) The computational cost of CoMoE. CoMoE leverages inactivated experts as negative samples, and its computational cost is proportional to $O(N)$. To alleviate the computational burden, a complexity reduction strategy and empirical validation are provided in Appendix H. However, it substantially improves expert specialization and modularization, achieving remarkable performance with a limited number of experts; (b) Other evaluation benchmarks were not considered. Nevertheless, CoMoE is simple and can be easily integrated with different backbone models and various downstream tasks.
@@
Acknowledgments
This work was supported in part by the Key Research Project of Chinese Academy of Sciences under Grant No. KGFZD-145-25-21, the National Natural Science Foundation of China under Grant 62322316, and the Open Fund of National Key Laboratory of Information Systems Engineering (No. 6142101240203).
@@ Appendix A Theoretical Proof
Proof. As illustrated in Eq. 8, we utilize $h_{1}(x,e^{+})\propto p(e^{+}|x)/p(e^{+})$ and $h_{2}(x,e^{-})\propto p(e^{-}|x)/p(e^{-})$ to approximate the information density ratio, which preserves the MI between $x$ and $M$. Based on the standard derivations presented in InfoNCE oord2018representation and wen2024contrastive, the contrastive objective can be rewritten as: